OpenStudy (anonymous):

Rextangle R has varying length L and width w with the constant perimeter of 4 feet How do I include a graph with relevant points labeled?

OpenStudy (anonymous):

make a square

OpenStudy (anonymous):

oh i see they want you to make a function that models this

OpenStudy (anonymous):

rectangle with length \(x\) and height \(y\) has area \(xy\) and perimeter \(2x+2y\) since the perimeter is \(4\) you have \(2x+2y=4\) making \(x+y=2\) or \(y=2-x\) that make the area \(A=xy\) the same as \[A(x)=x(2-x)\] or if you prefer \[A(x)=2x-2^2\] you can graph that one

OpenStudy (anonymous):

oops i mean if you prefer \[A(x)=2x-x^2\] that's better

OpenStudy (anonymous):

Would I just plug that equation in my calculator?